Select any two integers between -12 and +12 which will become solutions to a system of two equations.
Write two equations that have your two integers as solutions. Show how you built the equations using your integers. Solve your system of equations by the addition/subtraction method.

Answer by stanbon(75887)

Select any two integers between -12 and +12 which will become solutions to a system of two equations.
----
Example: Select x = -5 and y = +7
--------------------------------
Write two equations that have your two integers as solutions.
2x+3y = 11
-x+2y = 19
---------------------------

Show how you built the equations using your integers. Solve your system of equations by the addition/subtraction method.
----
Solve:
Multiply the 2nd equation by 2:
2x+3y = 11
-2x+4y = 38
-----
Add and solve for "y":
7y = 49
y = 7
-----
solve for "x":
2x+3y = 11
2x + 3*7 = 11
2x + 21 = 11
2x = -10
X = -5
=====================
cheers,
Stan H.